OpenStudy (anonymous):

the resultant of two forces 3N and 4N making an angle 90 with each other is a) 1N b) 7N c) 5N d) 3.5N

OpenStudy (abhisar):

Hello @needsknowledge ! You can add any two vectors using the vector addition law i.e \(\boxed{\sf A+B=\sqrt{{A^2+B^2+2ABcos\theta }}}\) Where, \(\theta\) is the angle betwenn two vectors.